40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4905     Entry time: Wed Jun 29 00:35:36 2011
Author: Koji 
Type: Update 
Category: LSC 
Subject: new LSC overview screen 80% done 

New LSC screen is 80% completed.

It is now accessible from the LSC menu of "sitemap".

Most of the part in the screen is clickable such that it launches another screen depending on the location of the click.

 

The bottom part of the screen still need some work.

RFPD screen is temporary

LSC control screen is also temporary

DAC overflow indicators are still broken.

Channel assignment of the whitening filters are arbitrary so far.

 

Attachment 1: LSC_OVERVIEW.png  46 kB  | Hide | Hide all
LSC_OVERVIEW.png
ELOG V3.1.3-